High-Low Cost Estimation and Profit Planning
Comparative 2007 and 2008 income statements for Dakota Products Inc. follow:

DAKOTA PRODUCTS INC.
Comparative Income Statements
For Years Ending December 31, 2007 and 2008
2007 2008
Unit sales 5,000 8,000
Sales revenue $75,000 $120,000
Expenses (82,000) (103,000)
Profit (loss) $(7,000) $17,000

(a) Determine the break-even point in units.
Answer

units

(b) Determine the unit sales volume required to earn a profit of $7,000.
Answer
Answer #1

Variable cost per unit = (103000-82000)/(8000-5000) = 7 per unit

Fixed cost = 103000-(8000*7) = 47000

Selling price per unit = 75000/5000 = 15

a) Break even unit = Fixed cost/Contribution margin per unit = 47000/(15-7) = 5875 Units

b) Required unit sales = (47000+7000)/8 = 6750 Units